a) A solution contains 0.34 M of a weak acid HA (Ka = 2.0 x 10-7) and 0.17 M NaA. What is the pH after 0.05 M of HCl is added to this solution (assume no volume change)

b) The pH of a solution containing 0.1 M of a weak acid HA is 6. Calculate Ka for this acid.

Note the acids are unrelated for the two problem parts.
Answer #1

a)

after 0.05 M HCl added

[HA] = 0.34 + 0.05 = 0.39 M

[NaA] = 0.17 - 0.05 = 0.12 M

pH = pKa + log [NaA] / [HA]

pKa = - log Ka = - log 2.0 x 10-7 = 6.70

pH = 6.70 + log [0.12]/[0.39]

pH = 6.20

b)

HA <---------> H+ + A-

initially

[HA] = 0.1 M

pH = 6.0

[H+] = 10-pH = 10-6.0 = 1.0 x 10-6 M

at equilibrium

[H+] = [A-] = 1.0 x 10-6 M

[HA] = 0.10 - 0.000001 = 0.099999

Ka = [H+][A-]/[HA]

Ka = [1.0 x 10-6][1.0 x 10-6] / [0.099999]

Ka = 1.0 x 10-11